OpenStudy (anonymous):

Margin error of sample mean A juice shop owner wants to know the average amount of juice his staff puts into each glass. A sample of 50 glasses was found to have an average (mean) of 245 ml. If the known standard deviation of the amount of juice in each glass is 4 ml for the entire population, what is the margin of error of the sample mean? A. 1.13 ml B. 1.24 ml C. 1.35 ml D. 1.49 ml

OpenStudy (anonymous):

I tried 4/sqrt(50) but that is none of the answers.

OpenStudy (perl):

ME= Zc * sigma / sqrt(n) ME = 1.96 * 4 / sqrt(40)

OpenStudy (anonymous):

how did you get sqrt(40)?

OpenStudy (perl):

typo

OpenStudy (perl):

ME = 1.96 * 4 / sqrt(50)= 1.108 at 95 % significance ME = 2.58 * 4 / sqrt(50)=1.459 at 99% significance

OpenStudy (anonymous):

I tried all the signifiance percents but none worked.

OpenStudy (anonymous):

I tried d first and it was wrong.

OpenStudy (anonymous):

I don't know how to choose between a b and c.

OpenStudy (perl):

im not sure, the significance level is not given. can you try to upload the question

OpenStudy (anonymous):

well I typed in the question exactly as written.

OpenStudy (perl):

oh they used Zc = 2

OpenStudy (perl):

ME = 2 * 4 / sqrt(50)

OpenStudy (perl):

that comes out exactly 1.13

OpenStudy (anonymous):

How did you get 2*4?

OpenStudy (perl):

whoever made this question must have thought this was clear or obvious (it isn't)

OpenStudy (perl):

$$ \bf \rm margin~ of~ error = ( critical ~z ~score ) * \frac{ standard~ deviation~ of pop.}{ \sqrt{sample size} } $$

OpenStudy (perl):

the critical z score is 2 here

OpenStudy (anonymous):

okay ill go with 1.13